OpenStudy (anonymous):

write the percent as a fraction or mixed number in lowest terms. 240%

OpenStudy (anonymous):

percent = out of 100 so 240% = 240/100 can you simplify that, @kharper7?

OpenStudy (anonymous):

240/100= 24/10= 12/5

OpenStudy (anonymous):

good (: